A stone is thrown vertically upward a with a velocity of 30m/s from a height of 50m. Two seconds later, a ball is thrown vertically upward with a velocity of 45m/s from a height of 35m. When and where will they be at the same level? What are their velocities at that point?


1
Expert's answer
2021-05-07T17:27:26-0400
"50+30t-0.5(9.8)t^2\\=35+45(t-2)-0.5(9.8)(t-2)^2\\t=3.6\ s"

"h=50+30(3.6)-0.5(9.8)(3.6)^2=94.5\ m"

"v_1=30-9.8(3.6)=-5.3\frac{m}{s}\\v_2=45-9.8(3.6-2)=29.3\frac{m}{s}"